# Transfer Learning on Edge Connecting Probability Estimation under Graphon Model

### v1

@@ C.2 Additional Simulation for Transfer Between Different Graphons
$6\rightarrow 8$ 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
$6\rightarrow 9$ 1.5 $\pm$ 0.4 1.6 $\pm$ 0.4 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
$6\rightarrow 10$ 2.3 $\pm$ 0.3 2.3 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
$7\rightarrow 1$ 2.0 $\pm$ 0.3 1.9 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
$7\rightarrow 2$ 1.3 $\pm$ 0.3 1.3 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
$7\rightarrow 3$ 1.6 $\pm$ 0.3 1.5 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
$7\rightarrow 4$ 1.3 $\pm$ 0.2 1.2 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
$7\rightarrow 5$ 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
$7\rightarrow 6$ 0.9 $\pm$ 0.2 1.0 $\pm$ 0.3 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
$7\rightarrow 9$ 1.7 $\pm$ 0.4 1.7 $\pm$ 0.4 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
$7\rightarrow 10$ 2.3 $\pm$ 0.4 2.2 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
$8\rightarrow 1$ 1.8 $\pm$ 0.3 1.5 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
$8\rightarrow 2$ 1.4 $\pm$ 0.3 1.4 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
$8\rightarrow 3$ 1.7 $\pm$ 0.3 1.6 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
$8\rightarrow 4$ 1.3 $\pm$ 0.2 1.3 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
$8\rightarrow 5$ 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
$8\rightarrow 6$ 1.0 $\pm$ 0.2 1.0 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
$8\rightarrow 9$ 1.9 $\pm$ 0.4 1.9 $\pm$ 0.3 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
$8\rightarrow 10$ 2.3 $\pm$ 0.3 2.3 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
$9\rightarrow 1$ 2.4 $\pm$ 0.3 2.3 $\pm$ 0.2 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
$9\rightarrow 2$ 1.5 $\pm$ 0.3 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
$9\rightarrow 3$ 1.9 $\pm$ 0.3 1.6 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
$9\rightarrow 4$ 1.6 $\pm$ 0.2 1.6 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
$9\rightarrow 5$ 1.6 $\pm$ 0.2 1.3 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
$9\rightarrow 6$ 1.2 $\pm$ 0.2 1.2 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
$9\rightarrow 7$ 1.9 $\pm$ 0.2 1.9 $\pm$ 0.3 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
$9\rightarrow 10$ 2.2 $\pm$ 0.3 2.3 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
$10\rightarrow 1$ 2.1 $\pm$ 0.3 2.0 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
$10\rightarrow 2$ 1.3 $\pm$ 0.3 1.2 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
$10\rightarrow 3$ 1.6 $\pm$ 0.3 1.5 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
$10\rightarrow 4$ 1.2 $\pm$ 0.2 1.1 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
$10\rightarrow 6$ 1.3 $\pm$ 0.4 1.4 $\pm$ 0.3 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
$10\rightarrow 7$ 1.9 $\pm$ 0.3 1.8 $\pm$ 0.3 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
$10\rightarrow 8$ 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
$10\rightarrow 9$ 1.5 $\pm$ 0.3 1.5 $\pm$ 0.4 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
C.3 Ablation Study
@@
Figure C2 presents the average estimation errors (MSE) of four configurations: GTrans (red diamond with solid line), GTrans-NonDebias (purple circle with dotted line), GTrans-NonSmooth (orange triangle with dashed line), and GTrans-Adj (yellow square with dash-dotted line). The x-axis represents the source sample size, and the y-axis shows the average MSE over 50 runs.
@@
Figure C2: Ablation study of different estimation methods: (a) illustrates the impact of increasing source sample size on estimation accuracy, and (b) demonstrates the effect of varying density shift ($\lambda$) on the performance of GTrans, GTrans-NonDebias, GTrans-NonSmooth, and GTrans-Adj.
C.4 Hyperparameter Selection
C.4.1 Debiasing Threshold $\delta$ Selection
@@
Figure C3: Heatmaps of the optimal threshold $\delta$ across different Graphon IDs. GTrans-GW (top) shows that $\delta=0.15$ is consistently optimal or near-optimal, while GTrans-EGW ($\epsilon=0.01$ (bottom) favors $\delta=0.18$ in most scenarios, with slight increases under high perturbation.
Figure C3 shows that the optimal threshold $\delta$ depends on the graphon’s inherent density and the perturbation level $\lambda$. For dense graphons, $\delta$ remains stable, reflecting robustness to structural shifts. In contrast, sparse graphons exhibit a sharp increase in $\delta$ under positive $\lambda$, as perturbations enhance connectivity and reduce noise sensitivity. The asymmetry arises from a bias–variance trade-off. Denser sources ($\lambda>0$) yield high-quality but mismatched transfer, favoring delayed debiasing with larger $\delta^{\ast}$; sparser sources ($\lambda<0$) require earlier correction, hence smaller $\delta^{\ast}$. Besides, GW distance increases with denser sources due to structural mismatch, and decreases with sparser ones despite degraded transfer—supporting a larger $\delta$ for dense, and smaller for sparse sources.
C.4.2 Regularization Parameter $\epsilon$ Selection
@@
To determine the optimal regularization parameter $\epsilon$ in EGW, we perform $K$-fold cross-validation over a candidate set $\epsilon_{\text{list}}=\{0.001,0.005,0.01,0.05,0.1\}$. This process evaluates the MSE between the estimated and true target graphon. We observe that $\epsilon=0.01$ is consistently selected as the best-performing choice across different settings. To further illustrate this, figure C4 presents boxplots of the average MSE for each $\epsilon$ under varying source sizes $100,200,300,\dots,1000$} with a fixed target size of $n_{t}=50$. Each boxplot shows the MSE distribution across ten graphon types. While some graphons like Graphon 9 benefit from larger $\epsilon$ (e.g., 0.1) due to oscillatory patterns, $\epsilon=0.01$ consistently achieves the lowest error, demonstrating its robustness across diverse structures.
Figure C4: Boxplot comparison of different regularization parameters $\epsilon$ for GTrans-EGW across varying source sample sizes. The results show that $\epsilon=0.01$ (red) consistently maintains lower variance and median MSE in most graphons, indicating its robustness.
Appendix D Additional Real Data Results
D.1 Additional results for graph augmentation task
In Transfer Learning, selecting the optimal source dataset and identifying the best class correspondence are crucial for effective knowledge transfer, as the structural characteristics of source labels significantly impact performance. To determine optimal class-to-class transfer pairings, we first estimate graphons separately for each source and target class using neighborhood smoothing [66]. Subsequently, we compute the pairwise Gromov–Wasserstein (GW) distances as well as the Entropic Gromov–Wasserstein (EGW) distances with a commonly chosen regularization parameter $\epsilon=0.01$ for each target-source graphon pair, identifying the best-matched class from the source domain for each target class based on structural similarity. We summarize the key structural statistics of all datasets used in our experiments in Table D3. Table D4 summarizes the GW distances calculated between each pair of source and target labels across the considered datasets, while Table D5 presents the corresponding EGW distances, reflecting the regularized alignment between graphon pairs.

### v2

@@ E.2 Additional Simulation for Transfer Between Different Graphons
6 $\rightarrow$ 7 1.8 $\pm$ 0.3 1.8 $\pm$ 0.3 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
6 $\rightarrow$ 8 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
6 $\rightarrow$ 9 1.5 $\pm$ 0.4 1.6 $\pm$ 0.4 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
6 $\rightarrow$ 10 2.3 $\pm$ 0.3 2.3 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
7 $\rightarrow$ 1 2.0 $\pm$ 0.3 1.9 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
7 $\rightarrow$ 2 1.3 $\pm$ 0.3 1.3 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
7 $\rightarrow$ 3 1.6 $\pm$ 0.3 1.5 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
7 $\rightarrow$ 4 1.3 $\pm$ 0.2 1.2 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
7 $\rightarrow$ 5 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
7 $\rightarrow$ 6 0.9 $\pm$ 0.2 1.0 $\pm$ 0.3 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
7 $\rightarrow$ 8 1.6 $\pm$ 0.2 1.6 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
7 $\rightarrow$ 9 1.7 $\pm$ 0.4 1.7 $\pm$ 0.4 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
7 $\rightarrow$ 10 2.3 $\pm$ 0.4 2.2 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
8 $\rightarrow$ 1 1.8 $\pm$ 0.3 1.5 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
8 $\rightarrow$ 2 1.4 $\pm$ 0.3 1.4 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
8 $\rightarrow$ 3 1.7 $\pm$ 0.3 1.6 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
8 $\rightarrow$ 4 1.3 $\pm$ 0.2 1.3 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
8 $\rightarrow$ 5 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
8 $\rightarrow$ 6 1.0 $\pm$ 0.2 1.0 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
8 $\rightarrow$ 7 1.3 $\pm$ 0.3 1.4 $\pm$ 0.3 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
8 $\rightarrow$ 9 1.9 $\pm$ 0.4 1.9 $\pm$ 0.3 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
8 $\rightarrow$ 10 2.3 $\pm$ 0.3 2.3 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
9 $\rightarrow$ 1 2.4 $\pm$ 0.3 2.3 $\pm$ 0.2 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
9 $\rightarrow$ 2 1.5 $\pm$ 0.3 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
9 $\rightarrow$ 3 1.9 $\pm$ 0.3 1.6 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
9 $\rightarrow$ 4 1.6 $\pm$ 0.2 1.6 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
9 $\rightarrow$ 5 1.6 $\pm$ 0.2 1.3 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
9 $\rightarrow$ 6 1.2 $\pm$ 0.2 1.2 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
9 $\rightarrow$ 7 1.9 $\pm$ 0.2 1.9 $\pm$ 0.3 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
9 $\rightarrow$ 8 2.0 $\pm$ 0.2 1.9 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
9 $\rightarrow$ 10 2.2 $\pm$ 0.3 2.3 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
10 $\rightarrow$ 1 2.1 $\pm$ 0.3 2.0 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
10 $\rightarrow$ 2 1.3 $\pm$ 0.3 1.2 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
10 $\rightarrow$ 3 1.6 $\pm$ 0.3 1.5 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
10 $\rightarrow$ 4 1.2 $\pm$ 0.2 1.1 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
10 $\rightarrow$ 5 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
10 $\rightarrow$ 6 1.3 $\pm$ 0.4 1.4 $\pm$ 0.3 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
10 $\rightarrow$ 7 1.9 $\pm$ 0.3 1.8 $\pm$ 0.3 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
10 $\rightarrow$ 8 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
10 $\rightarrow$ 9 1.5 $\pm$ 0.3 1.5 $\pm$ 0.4 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
E.3 Ablation Study
@@
Figure E2 presents the average estimation errors (MSE) of four configurations: GTrans (red diamond with solid line), GTrans-NonDebias (purple circle with dotted line), GTrans-NonSmooth (orange triangle with dashed line), and GTrans-Adj (yellow square with dash-dotted line). The x-axis represents the source sample size, and the y-axis shows the average MSE over 50 runs.
@@
Figure E2: Ablation study of different estimation methods: (a) illustrates the impact of increasing source sample size on estimation accuracy, and (b) demonstrates the effect of varying density shift ($\lambda$) on the performance of GTrans, GTrans-NonDebias, GTrans-NonSmooth, and GTrans-Adj.
E.4 Hyperparameter Selection
E.4.1 Debiasing Threshold $\delta$ Selection
@@
Figure E3: Heatmaps of the optimal threshold $\delta$ across different Graphon IDs. GTrans-GW (top) shows that $\delta=0.15$ is consistently optimal or near-optimal, while GTrans-EGW ($\epsilon=0.01$ (bottom) favors $\delta=0.18$ in most scenarios, with slight increases under high perturbation.
Figure E3 shows that the optimal threshold $\delta$ depends on the graphon’s inherent density and the perturbation level $\lambda$. For dense graphons, $\delta$ remains stable, reflecting robustness to structural shifts. In contrast, sparse graphons exhibit a sharp increase in $\delta$ under positive $\lambda$, as perturbations enhance connectivity and reduce noise sensitivity. The asymmetry arises from a bias–variance trade-off. Denser sources ($\lambda>0$) yield high-quality but mismatched transfer, favoring delayed debiasing with larger $\delta^{\ast}$; sparser sources ($\lambda<0$) require earlier correction, hence smaller $\delta^{\ast}$. Besides, GW distance increases with denser sources due to structural mismatch, and decreases with sparser ones despite degraded transfer—supporting a larger $\delta$ for dense, and smaller for sparse sources.
E.4.2 Regularization Parameter $\epsilon$ Selection
@@
To determine the optimal regularization parameter $\epsilon$ in EGW, we perform $K$-fold cross-validation over a candidate set $\epsilon_{\text{list}}=\{0.001,0.005,0.01,0.05,0.1\}$. This process evaluates the MSE between the estimated and true target graphon. We observe that $\epsilon=0.01$ is consistently selected as the best-performing choice across different settings. To further illustrate this, figure E4 presents boxplots of the average MSE for each $\epsilon$ under varying source sizes $100,200,300,\dots,1000$} with a fixed target size of $n_{t}=50$. Each boxplot shows the MSE distribution across ten graphon types. While some graphons like Graphon 9 benefit from larger $\epsilon$ (e.g., 0.1) due to oscillatory patterns, $\epsilon=0.01$ consistently achieves the lowest error, demonstrating its robustness across diverse structures.
Figure E4: Boxplot comparison of different regularization parameters $\epsilon$ for GTrans-EGW across varying source sample sizes. The results show that $\epsilon=0.01$ (red) consistently maintains lower variance and median MSE in most graphons, indicating its robustness.
Appendix F Additional Real Data Results
F.1 Additional results for graph augmentation task
In Transfer Learning, selecting the optimal source dataset and identifying the best class correspondence are crucial for effective knowledge transfer, as the structural characteristics of source labels significantly impact performance. To determine optimal class-to-class transfer pairings, we first estimate graphons separately for each source and target class using neighborhood smoothing [65]. Subsequently, we compute the pairwise Gromov–Wasserstein (GW) distances as well as the Entropic Gromov–Wasserstein (EGW) distances with a commonly chosen regularization parameter $\epsilon=0.01$ for each target-source graphon pair, identifying the best-matched class from the source domain for each target class based on structural similarity. We summarize the key structural statistics of all datasets used in our experiments in Table F3. Table F4 summarizes the GW distances calculated between each pair of source and target labels across the considered datasets, while Table F5 presents the corresponding EGW distances, reflecting the regularized alignment between graphon pairs.